You have $50,000 for down payment. Your monthly gross income is $5,000. Your monthly car loan payment is $300. You also need to pay $400 every month for property taxes and home insurance. The bank has a lending guideline for mortgage loan that total debt service (TDS) ratio cannot exceed 44% of monthly gross income. How much can you borrow from a 25-year conventional mortgage with an effective annual rate of 6%?
Answer #1

TDS ratio=(Mortgage payments+Property Taxes+Other Debt payments)/Gross Income=(400+300+Mortgage)/5000

=>(400+300+Mortgage)/5000<=44%

=>Mortgage<=1500

SO, mortgage payment per month is $1500

Hence, loan amount=PV((1+6%)^(1/12)-1,25*12,1500)=236361.60